What is the coefficient of static friction between a load of mass 0.75kg and a horizontal surface if the limiting frictional force is 5N? (assume g= 10m/s)
1
Expert's answer
2019-07-09T12:14:20-0400

The friction force

"F=\mu N=\mu mg"

So, the coefficient of static friction

"\mu=\frac{F}{mg}=\frac{5\:\rm{N}}{0.75\:\rm{kg}\times 10\:\rm{N/kg}}=0.67"